# Scoring, Reasoning, and Selecting the Best! Ensembling Large Language Models via a Peer-Review Process

### v1

@@ 2.3 Selecting the Best
GaC 71.5 91.8 54.0 23.6 60.2
LLM-PeerReview-Average 76.9 $\pm$ 0.1 92.7 $\pm$ 0.3 69.5 $\pm$ 0.2 30.4 $\pm$ 0.1 67.4
LLM-PeerReview-Weighted 77.0 $\pm$ 0.1 93.0 $\pm$ 0.2 71.0 $\pm$ 0.2 30.2 $\pm$ 0.1 67.8
Our variants Llama-3-8B-Selection 76.5 $\pm$ 0.2 90.8 $\pm$ 0.6 68.8 $\pm$ 0.5 29.6 $\pm$ 0.3 66.4
@@
Table 1: Main results (%).
Figure 3: LLM performances (bottom: AlpacaEval).
3 Experiments
3.1 Setup
We provide more details in Appendix A, including experimental setups, additional results, and the prompts used.
Datasets and evaluation.
We evaluate four widely-used datasets  [^4] We employ the same versions as those used in previous studies (Guha et al., 2024; Hu et al., 2024)., grouped into three categories: (1) Factual Recall: TriviaQA (Joshi et al., 2017; Guha et al., 2024) evaluates the accuracy of model responses to factual questions across various domains, including history, science, and geography. (2) Arithmetic Reasoning: GSM8k (Chen et al., 2021a; Hu et al., 2024) and MATH (Hendrycks et al., 2021; Hu et al., 2024) assess basic arithmetic and more advanced mathematical reasoning, respectively, with accuracy as the evaluation metric, focusing on correct numerical answers. (3) Instruction Following: AlpacaEval (Dubois et al., 2023; Hu et al., 2024) tests models’ ability to follow various instructions. We use GPT-4o-mini to evaluate the accuracy of model responses, assessing whether the model’s response exceeds the reference answer in the dataset.
Seed LLMs for ensemble.
Considering 7B-scale models are widely used by researchers and generally regarded as having acceptable judging capabilities (Wang et al., 2025; Kim et al., 2024), we use these well-established 7B models for ensemble: Llama-3.1-8B-Instruct, Mistral-7B-Instruct, Qwen2-7B-Instruct, and Qwen2.5-7B-Instruct. Furthermore, we employ the instruction-tuned versions of these models rather than their base counterparts. This choice allows us to leverage their instruction-following capabilities, which are essential for the LLM-as-a-judge tasks.
Baselines.
We compare the proposed LLM-PeerReview with the two categories of baselines. (1) Single LLMs: The four 7B-scale models, Llama-3.1-8B-Instruct, Mistral-7B-Instruct, Qwen2-7B-Instruct, and Qwen2.5-7B-Instruct. (2) LLM Ensemble baselines: (i) Random (Lu et al., 2024; Guha et al., 2024) is a random-selection baseline that simply returns the response from a randomly chosen LLM in the ensemble. As one of the simplest ensemble strategies for large language models, this method has previously been applied to dialogue tasks (Lu et al., 2024); (ii) Smoothie-Global (Guha et al., 2024), Smoothie-Local (Guha et al., 2024), and Agent-Forest (Li et al., 2024c) are recently proposed, strong similarity-based ensemble methods, as introduced in detail in Section 1; (iii) GaC (Yu et al., 2024) is a representative token-level ensemble-during-inference approach. It constructs a unified vocabulary that merges the individual dictionaries of multiple LLMs. During inference, token sampling is performed by observing the output distributions from these models across the unified vocabulary.
Configurations.
(1) For each individual large language model, we follow the setup of Smoothie (Guha et al., 2024), where the model responds once to each query. The responses from all models are stored for integration by the LLM Ensemble methods. (2) For the two variants of the baseline Smoothie (Guha et al., 2024), we set the number of neighbors as specified in the original paper. Agent-Forest (Li et al., 2024c) does not require any hyperparameter configuration. For our method, we set the model temperature to 0 during the scoring process to eliminate suboptimal results caused by randomness. Additionally, the scoring prompts used across the four datasets are provided in the Appendix. (3) All experiments were performed using 6 or 4 parallel Nvidia V100 32GB GPUs. All experiments with stochastic outputs were conducted three times.
(a) Left: The transition matrix of each LLM estimated by LLM-PeerReview-Weighted. Right: Correlation between matrix diagonal information of each LLM and its performance as a single judge (corresponding to “our variants” in Table 1). For the first three datasets with ground-truth answers, diagonal information is represented by extreme values $(\pi_{11}^{(j^{\prime})}+\pi_{KK}^{(j^{\prime})})$; for the instruction-following dataset AlpacaEval, the sum of all diagonal values is used.
Method TriviaQA GSM8k MATH AlpacaEval Average
Variant Performance ($\uparrow$):
   Random 65.7 $\pm$ 1.3 81.3 $\pm$ 0.6 51.2 $\pm$ 1.8 14.7 $\pm$ 0.5 53.2
   Single 69.2 $\pm$ 0.6 85.5 $\pm$ 2.1 60.3 $\pm$ 1.6 23.8 $\pm$ 1.0 59.7
   Double 73.3 $\pm$ 0.5 90.0 $\pm$ 0.7 71.3 $\pm$ 0.2 29.2 $\pm$ 0.2 66.0
   Flipped-triple 74.5 $\pm$ 0.0 90.8 $\pm$ 0.2 71.5 $\pm$ 0.4 30.5 $\pm$ 0.0 66.8
   Quadruple-half 74.7 $\pm$ 0.2 91.5 $\pm$ 0.4 73.3 $\pm$ 0.2 29.2 $\pm$ 0.2 67.2
Computation Efficiency (seconds/scoring the 4 responses for each sample; $\downarrow$):
   Single ($\mathcal{O}(J)$) 7.89 10.2 10.6 16.9 11.4
   Double ($\mathcal{O}(J^{2})$) 37.1 49.4 51.6 77.4 53.9
   Flipped-triple ($\mathcal{O}(J)$) 29.7 43.4 47.1 74.3 48.6
   Quadruple-half ($\mathcal{O}(J!)$) 51.3 83.8 90.0 137.65 90.7
Table 2: Top: Performance of the base variant LLM-PeerReview-Average with different scoring strategies. Note that, for computation efficiency, we use 200 samples from each dataset (all shuffled as necessary). (i) Random refers to the same baseline in Table 1; (ii) Single refers to scoring a single response at a time; (iii) Double refers to scoring all response pairs within the response set at a time; (iv) Quadruple-half refers to scoring all possible response quadruple sequences in the response set; given the high computational cost, we used a relaxed version and only calculated half of the cases. Bottom: Computation efficiency.
3.2 Main Results
The ensemble of the proposed LLM-PeerReview is effective. The main results are shown in Table 1. First, by examining the results in the “Single LLM” and “LLM Ensemble” sections of Table 1, one key finding is that both of our variants consistently outperform any single LLM and all LLM Ensemble baselines across all datasets. In the last column, which presents the average performance, our two variant methods (with results of 67.4% and 67.8%) surpass the strongest single model, Qwen2.5, by 4.7% and 5.1%, respectively, and outperform the strongest ensemble method, Smoothie-Global, by 6.9% and 7.3%. These results directly demonstrate the effectiveness of our method, as it achieves superior performance by integrating the collective knowledge of multiple models across factual-recall QA tasks, mathematical reasoning tasks, and instruction-following tasks. Additionally, the ensemble task across these four datasets is challenging, as the performance of the four LLMs varies significantly for each dataset. In contrast, ensembling four LLMs with similar performance would make it easier to achieve superior results compared to any single LLM.
Each LLM has its strengths and weaknesses. In Figure 3, the upper subplot presents a radar chart of individual LLM performance, while the lower subplot displays the win-tie-loss chart for models on the challenging instruction-following dataset, AlpacaEval. This chart highlights that models with the best overall performance may underperform on specific tasks compared to those with weaker overall results. In summary, the results in Table 1 and Figure 3 demonstrate that a strong LLM does not excel across all datasets. Each model has its strengths and weaknesses, highlighting the substantial practical significance of LLM Ensemble.
3.3 Significance of Aggregating Multiple Judges
Simply averaging the scores from multiple judges is also quite effective. In the “Our variants” of Table 1, we present the performance of using a single LLM as a judge to select the optimal response. From the average performances in the last column of Table 1, we observe that these variants perform quite well (surpassing the overall best model, Qwen2.5, in 3/4 cases). However, when comparing the performance of these variants with that of our prototype LLM-PeerReview-Average, it becomes clear that aggregating and averaging the scores from multiple judges is highly beneficial, compared to relying solely on the score of a single large model.
The weighted truth inference has the potential for further performance improvement. By observing the average results in Table 1, we find that LLM-PeerReview-Weighted leads to further performance gains compared to simple averaging. In the left subplot of Figure 11(a), we observe subtle variations in the transition matrices learned for each model. On the other hand, the right subplot in Figure 11(a), displaying positive correlation coefficients, demonstrates that our method can effectively identifies stronger and weaker judges.
3.4 Alternative Designs
The flipped-triple scoring trick represents a performance-efficiency trade-off. As introduced in Section 2.1, it is intuitive that, in addition to the our recommended flipped-triple scoring method, several variant scoring methods could be employed (their definitions are provided in the caption of Table 2). Overall, the performance of these four variants follows the order: $\textit{quadruple-half}>\textit{flipped-triple}>\textit{double}>\textit{single}$. Variants quadruple-half, flipped-triple, and double all offer noticeable de-biasing performance advantages over the single-scoring strategy. On the other hand, in terms of theoretical computational complexity, the complexities for de-biased strategies double/flipped-triple/quadruple-half are $\mathcal{O}(J)$$\mathcal{O}(J^{2})$$\mathcal{O}(J)$$\mathcal{O}(J!)$, with strategy flipped-triple having the lowest computational complexity. Furthermore, in Table 2, we present the scoring efficiency of these four strategies. Compared to strategies double and quadruple-half, strategy flipped-triple is the most time-efficient.
Common scoring levels can generally be attempted. In Figure 11, we conduct a further analysis of how different scoring levels influence the performance of both our method and the four individual scoring models. [^5] As indicated by the prompts in the Appendix, for the main experiment in Table 1, the scoring levels applied across the four datasets were 5, 3, 3, and 10, respectively. For each scoring level, we have carefully crafted meaningful descriptions and corresponding prompts. We use the basic variant LLM-PeerReview-Avarage for this analysis. Under these conditions, our method exhibits slightly varying performance, showing no consistent tendencies across the levels of 3, 5, 7, and 10.
4 Conclusion
Appendix A Appendix
  • Section A.2: More details on the experimental setup;
  • Section A.3: More experimental results;
  • Section A.4: The prompts used in the experiments.
A.1 Proof of the Optimization
The optimization objective is to maximize the log conditional likelihood of the observed scoring labels $\mathbf{Y}=\{y^{(i,j;j^{\prime})}\mid 1\leq i\leq I,\ 1\leq j\leq J,\ 1\leq j^{\prime}\leq J\}$ contributed by $J$ LLM judges, i.e., $\log p(\mathbf{Y};\Theta)$. In brief, as with most latent variable models (Bishop & Nasrabadi, 2006; Dawid & Skene, 1979), we apply the Expectation-Maximization (EM) algorithm (Dempter, 1977) to solve the optimization problem. First, the log-likelihood is:
$=$ $\sum_{i=1}^{I}\sum_{j=1}^{J}\log p(\mathbf{y}^{(i,j)};\Theta)$
used Jensen’s inequality (Bishop & Nasrabadi, 2006); and $q(t^{(i,j)})$ is a discrete distribution over the variable $t^{(i,j)}$. In the following, we then proceed to apply the general EM recipe to perform iterative calculations (concerning E-step and M-step) to solve the optimization problem: $\Theta:=\underset{\Theta}{\operatorname{argmax}}\log p(\mathbf{Y};\Theta)$.
[…]
M- step (learning): (A.5)
Furthermore, by maximizing optimization objective in Equation A.5 and using the standard Lagrange multiplier method (Bishop & Nasrabadi, 2006), we can obtain the closed-form solution for $\bm{\alpha}=\{\alpha^{(k)}\mid 1\leq k\leq K\}$ in Equation A.6 shown below; and by equating the gradient of Equation A.2 to zero, we can obtain the closed-form solution for $\{\bm{\Pi}^{(j^{\prime})}\}_{j^{\prime}=1}^{J}$ in Equation A.7 shown below.
$\alpha_{k}=\frac{\sum_{i=1}^{I}\sum_{j=1}^{J}q(t^{(i,j)}=k)}{I\cdot J},$ (A.6)
A.2 More Experiment Setup
The models used in the experiments are shown in Table n the following, we provide further descriptions of the four datasets used.
A.3 More Experimental Results
In this section, we provide further experimental results, including: more variant results, win-tie-loss charts for models across all four datasets, and four case studies, each corresponding to one dataset.
[…]
Single LLM Llama-3.1-8B-Instruct 75.3 79.3 52.3 7.3 53.5
Mistral-7B-Instruct 72.7 64.3 26.5 10.4 43.5
Qwen2-7B-Instruct 63.0 88.5 59.8 15.2 56.6
Qwen2.5-7B-Instruct 62.5 91.5 69.3 27.6 62.7
Theoretical average 68.4 80.9 51.9 15.1 54.1
Smoothie-Global 63.0 91.5 59.8 27.6 60.5
Smoothie-Local 73.6 85.5 61.8 18.3 59.8
Agent-Forest 70.5 86.8 61.0 22.1 60.1
LLM-PeerReview-Average 76.9 $\pm$ 0.1 92.7 $\pm$ 0.3 69.5 $\pm$ 0.2 30.4 $\pm$ 0.1 67.4
LLM-PeerReview-Weighted 77.0 $\pm$ 0.1 93.0 $\pm$ 0.2 71.0 $\pm$ 0.2 30.2 $\pm$ 0.1 67.8
Our variants (flipped-triple) Llama-3-8B-Selection 76.5 $\pm$ 0.2 90.8 $\pm$ 0.6 68.8 $\pm$ 0.5 29.6 $\pm$ 0.3 66.4
Our variants (single) Llama-3-8B-Selection 69.8 $\pm$ 0.3 83.7 $\pm$ 1.3 56.8 $\pm$ 0.4 21.5 $\pm$ 0.6 58.0
@@ A.3 More Experimental Results
Table A.2: Main results (%).
Figure A.1: Win-tie-loss charts on four datasets.
More variant results.
By examining Table A.2, we observe the following: (i) The four variants of “single” obtained through the “naive point-wise scoring” method outperform the “theoretical average” across all four datasets. This suggests that these variant methods have a distinct advantage over random selection from model responses, implying they are useful; (ii) However, these four variants are still insufficiently useful when compared to the flipped-triple variant methods in Table A.3, indicating a performance disadvantage. By contrasting the performance of “variants (single)” and “variants (flipped-triple)”, it becomes clear that our proposed flipped-triple scoring strategy is indeed effective. This conclusion supports the assertion that the flipped-triple strategy, as designed in Section 2.1 of the main paper, mitigates the bias introduced by the naive point-wise scoring method. Furthermore, it is precisely upon this superior “flipped-triple” scoring strategy that the performance of our two primary variants, LLM-PeerReview-Average and LLM-PeerReview-Weighted, exhibits its ultimate superiority.
Win-tie-loss charts.
@@
Case studies.
@@
Subsequently, in Figures A.2, A.3, A.4, and A.5, we present a case study for each corresponding dataset. Each case study illustrates the core computational process and final results of our method, as well as those of the comparison methods, Agent-Forest (Li et al., 2024c) and Smoothie (Guha et al., 2024), highlighting the key computational steps and outcomes. We provide a more detailed analysis of the first two datasets, as shown in the captions; the patterns observed in the remaining two datasets are consistent with those in the first two.
A.4 Prompts
@@
Figure A.2: Case study on dataset TriviaQA. Analysis: For this query, the correct answer is “Leander Club”. For each response generation, we followed the approach in  Guha et al. (2024) and applied truncation techniques under the same configuration. (i) For our method, the results of our variant, LLM-PeerReview-Average, are shown in the Figure; also, the selected Response after truth inference remains correct, with the four scalar values obtained as: [2.39, 2.39, 1.18, 1.18]; (ii) For the baseline Agent-Forest, it focuses on “which response has the highest cumulative BLEU similarity to all other responses”. Since Responses 1 and 2 are correct, while Responses 3 and 4 are incorrect, the cumulative BLEU similarity of Response 4 with all others is the highest, so Agent-Forest incorrectly selects Response 4; (iii) For the baseline Smoothie, its two variants do not consider the four responses for the current query during the calculation. Instead, they either focus on identifying the best overall model or the responses from the most similar neighboring query. As a result, both variants of Smoothie also make errors.
@@
Figure A.3: Case study on dataset GSM8K. The ellipses “…” in the Responses represent the omission of less important content. Analysis: The correct answer for this query is “2”. (i) For our method, the results of our variant, LLM-PeerReview-Average, are shown in the Figure; also, the selected Response after truth inference remains correct, with the four scalar values obtained as: [1.12, 1.00, 2.01, 2.90]; (ii) For the Agent-Forest method, it focuses on determining “which response has the highest cumulative BLEU similarity to all other responses”. Despite Response 2 not providing the correct final answer, Agent-Forest selects it because the weight (i.e., cumulative BLEU similarity score) is the highest. This is likely due to the high similarity between many words in the response compared to other responses. In this case, the simple analysis of “which response has the highest cumulative BLEU similarity” fails; (iii) For the Smoothie method, its two variants do not consider the four responses for the current query during the computation. Instead, they either focus on identifying the best overall model or consider the four responses of the most similar neighboring query. In contrast to the results shown in Figure A.2, here both variants of Smoothie obtained the correct result, despite the computation process being flawed.
@@
A prompt example (after query instantiation) for dataset MATH. Can you solve the following math problem? If $c$ is a constant such that $x^{2}+100x+c$ is equal to the square of a binomial, then what is $c$? Explain your reasoning. Your final answer should be a single numerical number, in the form ‘the answer is {answer}’, at the end of your response.

### v2

@@ A.5 More Experimental Results
LLM-PeerReview 76.9 $\pm$ 0.1 92.7 $\pm$ 0.3 69.5 $\pm$ 0.2 30.4 $\pm$ 0.1 67.4
LLM-PeerReview-W 77.0 $\pm$ 0.1 93.0 $\pm$ 0.2 71.0 $\pm$ 0.2 30.2 $\pm$ 0.1 67.8
Our variants (flipped-triple) Llama-3.1-8B-Selection 76.5 $\pm$ 0.2 90.8 $\pm$ 0.6 68.8 $\pm$ 0.5 29.6 $\pm$ 0.3 66.4
@@
Our variants (single) Llama-3.1-8B-Selection 69.8 $\pm$ 0.3 83.7 $\pm$ 1.3 56.8 $\pm$ 0.4 21.5 $\pm$ 0.6 58.0
@@
Table A.3: Main results.(%).
Figure A.1: Win-tie-loss charts for the 7B LLMs on four datasets.
Figure A.2: Comparison of key LLM Ensemble baselines versus LLM-PeerReview variants with different numbers of judges.
(a) More variant results.
By examining Table A.3, we observe the following: (i) The four variants of “single” obtained through the “naive point-wise scoring” method outperform the “theoretical average” across all four datasets. This suggests that these variant methods have a distinct advantage over random selection from model responses, implying they are useful; (ii) However, these four variants are still insufficiently useful when compared to the flipped-triple variant methods in Table A.3, indicating a performance disadvantage. By contrasting the performance of “variants (single)” and “variants (flipped-triple)”, it becomes clear that our proposed flipped-triple scoring strategy is indeed effective. This conclusion supports the assertion that the flipped-triple strategy, as designed in Section 2.1 of the main paper, mitigates the bias introduced by the naive point-wise scoring method. Furthermore, it is precisely upon this superior “flipped-triple” scoring strategy that the performance of our two primary variants, LLM-PeerReview and LLM-PeerReview-W, exhibits its ultimate superiority.
(b) Win-tie-loss charts.
@@
(c) Comparison of key LLM Ensemble baselines versus LLM-PeerReview variants with different numbers of judges.
The experimental results are illustrated in Figure A.2. As a supplementary evaluation to Section 4.3 in the main paper, we have the following findings: (i) For LLM-PeerReview, the average performance across four datasets (bottom sub-figure in Figure A.2) shows that increasing the number of judges generally improves performance (specifically, $our:1\text{ judge}<2\text{ judges}<4\text{ judges}=5\text{ judges}$). However, adding Llama-3-8B-Instruct  [^6] Huggingface URL: https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. as a fifth judge did not yield further gains, primarily because performance improvements become increasingly marginal as the ensemble size grows. Additionally, we found that Llama-3-8B-Instruct’s individual average performance (62.1%) is the lowest among the four candidate judges; for comparison, Llama-3.1-8B-Instruct achieved 66.4%; (ii) Furthermore, the overall average performance indicates that LLM-PeerReview significantly outperforms four classic LLM Ensemble baselines, including GaC, across all judge configurations. Regarding efficiency, as analyzed in Section 4.3, the variant using a single judge is more time-efficient than GaC.
(d) More experiments on 13B-scale LLMs.
In our main paper, we conducted experimental analysis using popular 7B-scale models. Here, we further investigate the performance of our method and baselines on 13B-scale LLMs. The corresponding experimental results and win-tie-loss charts for the LLMs are presented in Table A.4 and Figure  A.3. Our primary findings are as follows: (i) Similar to the 7B-scale results in Table 1 of the main paper, we obtained consistent findings for the 13B-scale models. The proposed methods, LLM-PeerReview and LLM-PeerReview-W, demonstrate significant advantages over both individual LLMs and established LLM Ensemble baselines; (ii) For the win-tie-loss charts, we observe a pattern similar to Figure 11: different models exhibit varying performance levels, manifesting dynamic and diverse strengths and weaknesses across different datasets.
[…]
Qwen2.5-14B-Instruct 73.5 34.5 54.0
DeepSeek-R1-Distill-Qwen-14B 59.5 17.5 38.5
Qwen2.5-14B-Instruct-1M 69.5 36.5 53.0
Phi-3-medium-4k-instruct 68.0 14.0 41.0
Theoretical average 67.6 25.6 46.6
Random 67.7 $\pm$ 0.6 26.0 $\pm$ 1.5 46.9
Smoothie-Global 73.5 34.5 54.0
Smoothie-Local 74.0 33.5 53.8
Agent-Forest 73.0 40.5 56.8
LLM-PeerReview 77.3 $\pm$ 0.2 46.5 $\pm$ 0.0 61.9
LLM-PeerReview-W 77.3 $\pm$ 0.2 47.3 $\pm$ 0.2 62.3
Table A.4: Main results for 13B-scale LLMs (%).
Figure A.3: Win-tie-loss charts for 13B-scale LLMs.
(e) Case studies.
@@
Subsequently, in Figures A.4, A.5, A.6, and A.7, we present a case study for each corresponding dataset. Each case study illustrates the core computational process and final results of our method, as well as those of the comparison methods, Agent-Forest (Li et al., 2024) and Smoothie (Guha et al., 2024), highlighting the key computational steps and outcomes. We provide a more detailed analysis of the first two datasets, as shown in the captions; the patterns observed in the remaining two datasets are consistent with those in the first two datasets.
@@
Figure A.4: Case study on dataset TriviaQA. Analysis: For this query, the correct answer is “Leander Club”. For each response generation, we followed the approach in  Guha et al. (2024) and applied truncation techniques under the same configuration. (i) For our method, the results of our variant, LLM-PeerReview, are shown in the Figure; also, the selected Response after truth inference remains correct, with the four scalar values obtained as: [2.39, 2.39, 1.18, 1.18]; (ii) For the baseline Agent-Forest, it focuses on “which response has the highest cumulative BLEU similarity to all other responses”. Since Responses 1 and 2 are correct, while Responses 3 and 4 are incorrect, the cumulative BLEU similarity of Response 4 with all others is the highest, so Agent-Forest incorrectly selects Response 4; (iii) For the baseline Smoothie, its two variants do not consider the four responses for the current query during the calculation. Instead, they either focus on identifying the best overall model or the responses from the most similar neighboring query. As a result, both variants of Smoothie also make errors.
@@
Figure A.5: Case study on dataset GSM8K. The ellipses “…” in the Responses represent the omission of less important content. Analysis: The correct answer for this query is “2”. (i) For our method, the results of our variant, LLM-PeerReview, are shown in the Figure; also, the selected Response after truth inference remains correct, with the four scalar values obtained as: [1.12, 1.00, 2.01, 2.90]; (ii) For the Agent-Forest method, it focuses on determining “which response has the highest cumulative BLEU similarity to all other responses”. Despite Response 2 not providing the correct final answer, Agent-Forest selects it because the weight (i.e., cumulative BLEU similarity score) is the highest. This is likely due to the high similarity between many words in the response compared to other responses. In this case, the simple analysis of “which response has the highest cumulative BLEU similarity” fails; (iii) For the Smoothie method, its two variants do not consider the four responses for the current query during the computation. Instead, they either focus on identifying the best overall model or consider the four responses of the most similar neighboring query. In contrast to the results shown in Figure A.2, here both variants of Smoothie obtained the correct result, despite the computation process being flawed.
@@ A.6 Prompts
A prompt example (after query instantiation) for dataset MATH. Can you solve the following math problem? If $x$ is a constant such that $x^{2}+100x+c$ is equal to the square of a binomial, then what is $c$? Explain your reasoning. Your final answer should be a single numerical number, in the form ‘the answer is {answer}’, at the end of your response.